It says to solve using the addition method. ???
-2x+y=-6
2x-3y=-2

Thnx

Answer by user_dude2008(1862)

-2x+y=-6
2x-3y=-2

-->

(-2x+2x)+(y-3y)=-6-2

-2y=-8

y=4

---->

-2x+4=-6

-2x=-10

x=5

Ans: x=5, y=4